OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (x+4)^2/9 - (x-5)^2/16 = 1

OpenStudy (goformit100):

For solving this do you know the conception of : Polynomial Function

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

@goformit100

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

That is the equation of a hyperbola whose transverse axis is horizontal, center (h,k), foci c units right and c units left of the center.

OpenStudy (anonymous):

vertices: (5, -1), (5, -7); Foci: (5, -9), (5, 1) @Mertsj

OpenStudy (mertsj):

You should put the x coordinate first in your ordered pairs.

OpenStudy (anonymous):

sooo its Vertices: (-1, 5), (-7, 5); Foci: (-9, 5), (1, 5) @Mertsj

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thnx

OpenStudy (mertsj):

yw